# Breaking the Dimensional Barrier: A Pontryagin-Guided Direct Policy Optimization for Continuous-Time Multi-Asset Portfolio Choice

### v5

@@ D.4 Deep BSDE Implementation Details
  • Policy Derivation: The optimal portfolio $\bm{\pi}_{t}$ is not a direct network output but is dynamically constructed at each step. For CRRA utility, the FOC-derived policy simplifies significantly. The simplification stems from the value function’s known separable form in wealth (i.e., $V\propto X_{t}^{1-\gamma}$). This property allows the general intertemporal hedging demand term, $-\frac{1}{X_{t}V_{xx}}\Sigma^{-1}(\bm{\sigma\rho\sigma_{Y}})V_{x\mathbf{y}}$, to be analytically transformed into the more tractable expression $\frac{1}{\gamma}\Sigma^{-1}\left(\bm{\sigma}\bm{\rho}\right)\left(\bm{\sigma}_{Y}\frac{\nabla_{\mathbf{y}}V}{V}\right)$.
@@
  • Initial State Sampling: For each training path, the initial time $t_{0}$ is sampled uniformly from the discrete time steps over $[0,T-\Delta t]$. The initial wealth $X_{0}$ is drawn uniformly from $[0.1,3.0]$, and initial factor states $\mathbf{Y}_{0}$ are sampled uniformly from a hyperrectangle representing their typical dynamic range (e.g., based on ±$\pm 3$ stationary standard deviations around their long-term means).

### v6

@@ E.5 Deep BSDE Implementation Details
  • Policy Derivation: The optimal portfolio $\bm{\pi}_{t}$ is not a direct network output but is dynamically constructed at each step. For CRRA utility, the FOC-derived policy simplifies significantly. The simplification stems from the value function’s known separable form in wealth (i.e., $V\propto X_{t}^{1-\gamma}$). This property allows the general intertemporal hedging demand term, $\frac{1}{X_{t}V_{xx}}\Sigma^{-1}(\bm{\sigma\rho\sigma_{Y}})V_{x\mathbf{y}}$, to be analytically transformed into the more tractable expression $\frac{1}{\gamma}\Sigma^{-1}\left(\bm{\sigma}\bm{\rho}\right)\left(\bm{\sigma}_{Y}\frac{\nabla_{\mathbf{y}}V}{V}\right)$.
@@
  • Initial State Sampling: For each training path, the initial time $t_{0}$ is sampled uniformly from the discrete time steps over $0,T-\Delta t]$. The initial wealth $X_{0}$ is drawn uniformly from $[0.1,3.0]$, and initial factor states $\mathbf{Y}_{0}$ are sampled uniformly from a hyperrectangle representing their typical dynamic range (e.g., based on $\pm 3$ stationary standard deviations around their long-term means).
@@
Appendix F Additional Numerical Results - Convergence Speed, Accuracy, and Policy Error Analysis
(a) n=1,k=10n=1,\;k=10
(b) n=10,k=10n=10,\;k=10
(c) n=50,k=10n=50,\;k=10
Figure 5: Average policy RMSE comparison between Baseline PG-DPO (e.g., blue), Projected PG-DPO (P-PGDPO) (e.g., orange), and Deep BSDE (e.g., green) relative to the benchmark solution across different asset dimensions ($n$) with $k=10$ factors. Values averaged over $k=10$ evaluation ‘slices.’ For each slice $j=1,\dots,10$, the evaluation (e.g., for RMSE calculation or visualization) focuses on state factor $Y_{j}$, while other factors $Y_{l},l\neq j,$ are held constant at their respective long-term means $\theta_{Y,l}$. Note the logarithmic scale on the $y$-axis.
Figure 5 shows the average Root Mean Squared Error (RMSE) between the learned or derived policies and the semi-analytical benchmark policy. The RMSE is plotted against the number of training iterations on a logarithmic scale for experiments with $k=10$ state variables and varying asset counts ($n=1,10,50$). For P-PGDPO, the reported RMSE corresponds to the total portfolio policy $\bm{\pi}^{\mathrm{PMP}}$ (Eq. (3.2)), which combines the myopic and intertemporal hedging components. As discussed in Section 5.2, the myopic error is typically negligible—on the order of $O(10^{-7})$. Thus, the total RMSE for P-PGDPO is effectively governed by the accuracy of the hedging component, which depends on estimating the costate gradient $\partial_{\mathbf{Y}}\lambda_{t}^{*}$ via BPTT.
Figures 6 and 7 display the error maps across models. In particular, Figure 6 illustrates the policy errors for the first asset ($i=1$) across three models-baseline PG-DPO, P-PGDPO, and Deep BSDE- in the high-dimensional setting with $n=50$ risky assets and $k=10$ state variables. The figure plots the difference between the total portfolio policy (i.e., the sum of the myopic and intertemporal hedging demands) generated by each method and the corresponding analytic benchmark solution. Panel (a) shows the Baseline PG-DPO, where substantial deviations from the benchmark are visible, particularly near maturity and for extreme values of $Y_{1}$. Panel (b) reports the results for P-PGDPO, which achieves significantly lower errors overall. The projection step based on stabilized costates produces policies that are visually close to the benchmark, with errors an order of magnitude smaller than in the baseline. Panel (c) presents the Deep BSDE method, which captures the broad structure of the policy but exhibits notable inaccuracies in regions where intertemporal hedging demand is critical.
(a) Baseline PG-DPO Error
(b) P-PGDPO Error
(c) Deep BSDE Error
Figure 6: Policy errors for the first asset ($i=1$) across models in the case where $n=50,k=10$. Errors are plotted against $Y_{1}$ at a representative time-to-maturity TtT-t (e.g., $t=(T-T_{0})/2$ if $T_{0}$ is initial time, assuming $T=1.5$ is the horizon from $t=0$; thus, for instance, at $t=0.75$ or time-to-maturity $0.75$). Other state factors $Y_{j\neq 1}$ are held at their long-term means $\theta_{Y,j}$.
(a) Myopic Error
(b) Hedging Error
Figure 7: Decomposition errors of the Projected PG-DPO (P-PGDPO) policy for the first asset ($i=1$) in the $n=50,k=10$ case (results from iteration 6,400 as plotted), plotted against $Y_{1}$ at a representative time-to-maturity TtT-t. The first figure shows the error for myopic demand and the second figure shows the error of intertemporal hedging demand. Other state factors $Y_{j\neq 1}$ are held at their long-term means $\theta_{Y,j}$.
In addition to the total portfolio errors, Figure 7 provides a decomposition of the policy errors into the myopic and intertemporal hedging components for the first asset ($i=1$) under the setting with $n=50$ risky assets and $k=10$ state variables. The left panel shows the error map for the myopic demand, while the right panel shows the error for the intertemporal hedging demand, both plotted against the first variable $Y_{1}$ and time-to-maturity (Tt)(T-t), with all other state variables fixed as their long-run means. The results demonstrate that the myopic demand is recovered with residual errors on the order of $10^{-6}$, confirming the robustness of the costate-driven projection step for the static component of the portfolio. By contrast, the hedging demand exhibits larger and more structured errors, concentrated in regions where horizon effects are most pronounced. This reflects the intrinsic difficulty of estimating the cross-derivatives of the value function that drive intertemporal hedging behavior, particularly in high-dimensional environments.
In sum, Figures 6 and 7 illustrate that while P-PGDPO achieves highly accurate recovery of both myopic and hedging components, the primary source of residual policy error originates from the hedging term, underscoring the importance of our methodological focus on stabilizing and projecting costate estimates to capture horizon-dependent dynamics.